OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (freckles):

tried to find the third term in the sequence by replacing n with 3

OpenStudy (freckles):

\[f(n)=-4 \cdot f(n-1)-3 \cdot f(n-2) \\ \] I see three n's to replace here with 3 \[f(3)=-4 \cdot f(3-1)-3 \cdot f(3-2) \\ \] now finish the operations inside the ( )

OpenStudy (freckles):

\[f(3)=-4 \cdot f(2)-3 \cdot f(1)\] as you see we need f(2) and f(2) is actually the 3rd term since f(0) is the first term

OpenStudy (freckles):

so let's find the the third term f(2) then we will go back to finding the fourth term f(3) so replace all the n's with 2's

OpenStudy (freckles):

\[f(n)=-4 \cdot f(n-1)-3 \cdot f(n-2) \\ \] go ahead try replace the n's you see there with 2's and do nothing else for now

OpenStudy (anonymous):

I don't know where to start

OpenStudy (freckles):

i'm telling you exactly where to start

OpenStudy (freckles):

you should see 3 n's replace the n's there with 2's

OpenStudy (freckles):

well there was one more n

OpenStudy (anonymous):

f(2) = -4 * f(2-1) -3 * f(n-2) ?

OpenStudy (freckles):

well there is that last n there you still haven't replaced with 2 \[f(2)=-4 \cdot f(2-1)-3 \cdot f(2-2)\]

OpenStudy (freckles):

now do the operations in the f( )

OpenStudy (freckles):

2-1=? 2-2?

OpenStudy (anonymous):

2-1 = 1 2-2 = 0

OpenStudy (freckles):

right so we have \[f(2)=-4 \cdot f(1)-3 \cdot f(0)\]

OpenStudy (freckles):

but guess what

OpenStudy (freckles):

f(1) and f(0) was given

OpenStudy (freckles):

remember this: f(0) = -2 f(1) = 8

OpenStudy (freckles):

replace f(1) with 8 replace f(0) with -2

OpenStudy (freckles):

\[f(2)=-4 \cdot f(1)-3 \cdot f(0) \\ f(2)=-4 \cdot 8-3 \cdot -2\] when folloing the order of operations we know we are to take care of the multiplication then the subtraction so can you simplify this

OpenStudy (anonymous):

-4 * 8 - 3 * -2? -26

OpenStudy (freckles):

-32+6=-26 sounds good

OpenStudy (freckles):

so far we have the sequence is f(0),f(1),f(2),f(3),f(4),.... where we have -2,8,-26,........

OpenStudy (freckles):

there are two choices you have that fit this so far

OpenStudy (freckles):

so we need to make another round

OpenStudy (freckles):

now let's go back to our f(3)

OpenStudy (freckles):

we already did some simplification but let's start from the top to find the 4th term (or the 3-1th term) replace the n's with 3 \[f(n)=-4 \cdot f(n-1)-3 \cdot f(n-2) \\ \] whereever you n put a 3 \[f(3)=-4 \cdot f(3-1)-3 \cdot f(3-2) \\ \]

OpenStudy (freckles):

see of you can simplify that you have to use previous terms in your sequence to finish simplifying

OpenStudy (freckles):

see if you can*

OpenStudy (anonymous):

what about the f's?

OpenStudy (anonymous):

3-1 is 2 3-2 is 1

OpenStudy (freckles):

you remember before we used what were given about the f's we were given f(0) and f(1)... here you have f(2) and f(1) we just found f(2) and we were given f(1)

OpenStudy (freckles):

right so you have \[f(3)=-4 \cdot f(2)-3 \cdot f(1)\] we just found f(2) to be -26 and we were given f(1) to be 8

OpenStudy (anonymous):

so is that f(3) = -4 * -26 - 3 * 8

OpenStudy (freckles):

beautiful

OpenStudy (anonymous):

80. Answer is B

OpenStudy (freckles):

cool stuff :)

OpenStudy (freckles):

you did well it just takes a bit of experience with these to get used to them

OpenStudy (anonymous):

thank you :)